# Understanding Iterative Combinatorial Auction Designs via Multi-Agent Reinforcement Learning

### v1

@@ Abstract.
Iterative combinatorial auctions are widely used in high stakes settings such as spectrum auctions. Such auctions can be hard to understand analytically, making it difficult for bidders to determine how to behave and for designers to optimize auction rules to ensure desirable outcomes such as high revenue or welfare. In this paper, we investigate whether multi-agent reinforcement learning (MARL) algorithms can be used to understand iterative combinatorial auctions, given that these algorithms have recently shown empirical success in several other domains. We find that MARL can indeed benefit auction analysis, but that deploying it effectively is nontrivial. We begin by describing modelling decisions that keep the resulting game tractable without sacrificing important features such as imperfect information or asymmetry between bidders. We also discuss how to navigate pitfalls of various MARL algorithms, how to overcome challenges in verifying convergence, and how to generate and interpret multiple equilibria. We illustrate the promise of our resulting approach by using it to evaluate a specific rule change to a clock auction, finding substantially different auction outcomes due to complex changes in bidders’ behavior.
conference: ; ;
@@ 1. Introduction
Consider the case of spectrum auctions. The iterative combinatorial auctions used to sell radio spectrum can have 50+ page manuals explaining how they evolve over time and which bids are legal to submit (FCC, 2015). They are held too infrequently to learn about from historical data, occurring only once ever few years nationally. What’s more, while there are a few popular auction formats, such as Simultaneous Multiple Round Auctions (SMRA) (Milgrom, 2000), Combinatorial Clock Auctions (CCA) (Ausubel et al., 2006), and Clock Auctions [^1] Clock Auctions are combinatorial auctions, but are a distinct auction format from Combinatorial Clock Auctions. (Ausubel, 2004), even within a single format, rules tend to differ from one auction to the next. For example, between the (consecutive) Canadian 700 MHz (ISED, 2012) and 600 MHz (ISED, 2018) auctions, the activity rule was altered; later, between the Canadian 3500 MHz (ISED, 2021) and 3800 MHz (ISED, 2022) auctions, a spectrum cap was added limiting the maximum a bidder could win in each region. Given that rules are in constant flux, and that it is difficult to predict how isolated rule changes will impact an auction’s equilibrium, it is important to develop methods for reasoning about the effects of counterfactual policy changes in these auctions.
@@
Ideally, we would analyze such problems by computing the Bayes–Nash equilibria of the game induced by each set of auction rules, evaluating how outcomes change in equilibrium. This is the dominant approach taken by theoretical work in economics, and has also been adopted by various computational approaches (e.g. Thompson et al., 2017; Thompson and Leyton-Brown, 2017; Rabinovich et al., 2013). However, there is little hope of identifying Bayes–Nash equilibria of realistic iterative combinatorial auctions, as they are too complex to admit pen-and-paper analysis using known techniques. Thus, the most influential existing work studies highly simplified settings. For instance, for one auction format, Riedel and Wolfstetter (2006) proved that it is an equilibrium for bidders to end the auction in the first round, but their proof applies only to auctions with a single product, infinitesimal price increases, and complete information about bidders’ values. Combinatorial action spaces and multi-round structures imply enormous extensive-form representations, putting them out of reach of traditional equilibrium solvers as well.
@@
An alternative is to simulate the auction under the assumption that bidders play some given, fixed strategy (e.g., Newman et al., 2024; Cary et al., 2007). One can then observe both the effect of a rule change and its sensitivity to changes in bidders’ valuations and to randomness in the auction mechanism. However, combinatorial auctions are not typically strategyproof (see e.g., Bosshard and Seuken, 2021; Levin and Skrzypacz, 2016). There is thus typically not a singular reasonable strategy to simulate, let alone one with consistent incentive properties across the rules being studied.
@@
Multi-agent reinforcement learning (MARL) offers a middle ground: agents who exhibit much richer strategic reasoning than in static simulations but much more computational tractability than classical Bayes–Nash equilibrium computation. Over the past decade, MARL algorithms have made great strides. Early efforts focused on achieving superhuman performance in two-player zero-sum games, such as Go (Silver et al., 2016), Shogi (Silver et al., 2017), and Stratego (Pérolat et al., 2022). These games share the special property that players do not need to solve the equilibrium selection problem (Harsanyi and Selten, 1988): if a player plays their part of an equilibrium strategy, it will do at least as well against any other strategy an opponent might play. Of course, this property does not hold in auctions, nor does it hold in zero-sum games having more than two players. It is thus striking that recent work has shown superhuman performance in multiplayer zero-sum games such as poker (Brown and Sandholm, 2019b). Of particular relevance to what follows, Counterfactual Regret Minimization (CFR) (Zinkevich et al., 2007) is a popular, modern MARL algorithm that was recently used to solve Texas Hold ’em poker (Bowling et al., 2015; Brown and Sandholm, 2018). Another impressive recent achievement was the demonstration of human-level play in the board game Diplomacy (Anthony et al., 2022; Bakhtin et al., 2022). Indeed, this game shares other key similarities with our domain: a combinatorial action space and iterative, simultaneous moves.
@@
We foresee no near-term interest from market participants in deploying “superhuman” bots to bid autonomously in complex and high-stakes auctions without human oversight—nobody is going to bet a Fortune 500 company’s future on an RL algorithm running against unknown opponent strategies in a novel economic environment. However, even in such domains, MARL tools offer promise for evaluating competing auction designs and gaining economic understanding. In this vein, MARL has been used to identify and evaluate complex policy decisions regarding taxation (Zheng et al., 2022) and regulations on economic platforms (Wang et al., 2023). MARL has also been applied to bargaining, where prior work has studied how different learning algorithms lead to more or less socially desirable outcomes (Abramowicz, 2020; Li et al., 2023). We are aware of fewer papers that use RL to study auctions specifically. Banchio and Skrzypacz (2022) study the behaviour of $Q$-learners in first- and second-price auctions, arguing that in fast-paced online advertising auctions such algorithms really do interact with each other and hence it is more important to understand their joint dynamics than to consider equilibrium behavior. Deep MARL has also been used to compute equilibria of single-round combinatorial auctions (Bichler et al., 2021, 2023). Closest to the domain we study, Pacaud et al. (2023) combine heuristic strategies with Monte-Carlo tree search to find effective bidding strategies in clock auctions; however, they focus on finding strategies that perform well against existing heuristics in auctions with complete information.
@@
This paper aims to use MARL to gain economic insight in (incomplete-information) iterative combinatorial auctions. Such insights could help a human bidder assemble a strategic “playbook” by providing examples of strong bidding both by an agent with their own preferences and by their counterparties; to date, such playbooks are typically derived via pen-and-paper analysis of dramatically simplified subproblems and via observations of expert human play in a handful of “mock auctions”. Insights gleaned from MARL analysis could also help an auction designer to trade off the costs and benefits of candidate rule changes, evaluating economic variables such as revenue, welfare, length of the auction, variance in outcomes, etc. Unfortunately, actually running MARL to understand an iterative combinatorial auction is not nearly as simple as implementing an auction simulator and then unleashing multiple copies of some off-the-shelf algorithm. Modeling a spectrum auction verbatim from its rulebook produces games that are vastly too large to solve. One is then immediately faced with important modeling choices, finding ways to reduce the game to a feasible size without abstracting away the most strategically important elements.
@@
Our work makes two main contributions. First, in Section 2, we present a general methodology for studying complex, iterative auctions using MARL. Notably, we advocate for simplifying the auction model by abstracting the action space, controlling the number of bidders, and limiting the auction’s length, without giving up on qualitatively important features such as imperfect information or asymmetry between bidders. We describe the landscape of MARL algorithms and important problems that can arise, such as arbitrary behavior when multiple actions lead to the same outcome, and convergence to brittle equilibria that involve unrealistic levels of coordination. We also discuss key challenges in verifying an algorithm’s convergence, and interpreting trained policies in the face of multiple equilibria.
@@
Second, we present a case study showing that, despite the challenges just described, MARL can be used to conduct novel, economically meaningful analysis. We consider a prominent family of iterative combinatorial auctions called clock auctions and, leveraging two modern MARL algorithms, investigate a key design choice: how to process bids when multiple bidders want to drop the same product to a level below supply. Section 3 describes the problem, how we model it as a game, how we solved it using MARL, and how we assessed performance. Section 4 then goes on to demonstrate that this rule change does indeed give bidders incentive to change their behavior, yielding substantial differences in both revenue and auction length. Furthermore, we show that modeling bidders as following a straightforward, myopic bidding heuristic gives misleading results in the same setting, and indeed would lead a designer to conclude that the rule change has the opposite effect.
@@ 2. Methodology: Understanding Iterative Combinatorial Auctions via Multi-Agent Reinforcement Learning
We begin by laying out our methodology for applying MARL to understand iterative, combinatorial auctions. First, we discuss how we model the auction, trading off model fidelity with size of the resulting game tree. We then turn to “solving” the game using MARL, discussing the strengths and weaknesses of different MARL algorithm families, how to address the problem of multiple equilibria, and considering equilibrium refinements (preferring pure equilibria; avoiding brittle equilibria). The section concludes by considering how to validate and interpret policies: assessing the extent to which policies have converged to a Bayes–Nash equilibrium and reasoning across multiple equilibria.
@@ 2.1. Modelling an Auction
It is tempting to simply model an auction wholesale, mapping its entire rules manual into a MARL environment, and then unleashing a MARL algorithm. Unfortunately, this approach is very unlikely to succeed for iterative combinatorial auction problems because the number of bundles will be large, the number of bidders will be large, and the auction will last many rounds, yielding an enormous game tree that is simply infeasible to solve. Much like pen-and-paper analysis, one must abstract away certain elements to buy the ability to accurately model others.
@@
We recommend gaining a solid foundation by starting with simple, easy-to-solve games and scaling up later rather than starting with complex ones that might prove to be intractable. We have been surprised countless times by how even auctions with moderate numbers of bidders, goods, and rounds can lead to an enormous, unwieldy game tree, leaving little hope of obtaining reliable results from MARL algorithms. However, as we demonstrate in our experiments in Section 3, there is a sweet spot of game sizes out of reach of pen-and-paper analyses, but very much tractable for MARL. We begin by describing several high-level decisions that a designer can make to control the size of the game tree. The key metric of importance is usually the number of distinct information states in the game—the number of game states that an agent can distinguish.
@@
Combinatorial auctions typically have enormous action spaces, often corresponding to the set of all legal bundles. The action space determines the branching factor of the game tree, so reining it in has a dramatic effect on game size. One helpful way to limit the size of an action space is to experiment with auctions that have few goods or units for each good available, limiting the number of possible bids. However, we caution against reducing the auction down to a single product (e.g., several units of a generic license), as auctions with only one product may not to engage with many of the auction’s rules, leaving interesting behavior unstudied. We recommend beginning with auctions offering two or three products.
@@
It is also possible to study auctions with more available goods by restricting bidders more heavily, only allowing them to choose from a small number of heuristic strategies at each state. For example, these heuristics could include bidding myopically, maintaining the last round’s bid, raising prices on an opponent’s holdings, or attempting to drop out of the auction; of course, many other options are possible. Care is required here to ensure that this restriction does not significantly hamper players’ ability to best respond to others’ strategies. As a concrete example, one might reduce the action space by limiting bidders to only bidding on packages they genuinely value. However, it is well-known that there are auction instances where such simple bidding is never optimal in Bayes–Nash equilibrium (Bosshard and Seuken, 2021). A helpful intermediate option is to allow players to deviate from a given set of heuristics up to $k$ times per auction; this can be useful when it is only strategically important to deviate from known heuristics occasionally, and also can be used to diagnose the effectiveness of the given set of heuristics. This approach has the beneficial side effect of producing strategies that are easier for an analyst to interpret, which can be particularly helpful for developing a playbook. Restricting actions to a smaller set is also likely to make the resulting strategies more useful as meta-strategies for use in Empirical Game Theory Analysis (Wellman, 2006).
@@
In some auctions, a naive implementation would produce a continuous action space. For instance, both intra-round bidding in clock auctions and the supplementary rounds of CCAs give bidders the ability to report real-valued prices. While there is some work on RL with continuous actions, we recommend against taking this route: these methods typically make strong assumptions that are not suitable in auction settings (e.g., assuming that bids are drawn from a Gaussian distribution). The simplest method, which we adopt in this work, is to simply avoid modelling the part of the auction that admits a continuous action space. When this is not possible, a good alternative is to restrict possible bids to a fixed grid: for instance, allowing bidders to only bid in \$1 increments. We suggest experimenting with multiple discretizations that vary in granularity to ensure that the chosen discretization does not have a substantial effect on the results. Finally, one can again employ a finite set of heuristics (each of which maps to an unrestricted real value), recovering a discrete action space without discretizing bids themselves.
@@
A related technical issue is that, in theory, auctions can be arbitrarily long: as long as bidders continue to overdemand products, the auction will continue. These arbitrarily large game trees do not typically pose a problem to theorists, who can handle them with inductive proofs, nor to reasonable bidders in practice, who would never expose themselves to indefinitely rising prices. However, they do pose a problem to most RL algorithms, which cannot learn to avoid each of these unreasonable bid trajectories without first gaining experience by making them. We recommend constraining agents’ bids to rule out such arbitrary-length bidding sequences altogether. A natural solution is to disallow bidders from making bids that are strictly dominated by dropping out of the auction, ensuring that they would eventually be forced to drop out when prices become sufficiently high. Another is modelling bidders as budget-constrained, disallowing bids that would exceed a bidder’s budget if processed. Budgets are a more tenuous modelling choice: while real-world bidders do truly have finite resources, it may be unrealistic to prohibit a bidder from buying a license that they value more than its cost. A further concern is that making an over-budget bid does not mean that the bidder will necessarily end up paying more than their budget; preventing such moves may unwittingly restrict the set of equilibria. We also experimented with implementing a fixed limit on auction length, giving all players arbitrary, negative utilities upon hitting this limit; we recommend against this as it can have a dramatic effect on the equilibria of the game. [^2] Specifically, we found that weak bidders that would otherwise be allocated nothing could make non-credible threats to bid up to the limit, forcing stronger bidders to concede an item.
@@
In practice, spectrum auctions can involve dozens of bidders. However, many bidders in real auctions represent small, locally-operating telecoms that have limited ability to make large-scale bids, and relatively few bidders make strategic choices at a national scale. If smaller bidders need to be modeled, consider whether they need the flexibility of strategic adaptation, whether their action spaces can be restricted to their bundles of interest, or whether a deterministic heuristic strategy would suffice.
@@
Real auctions, where bidders likely have priors over their opponents’ strategic goals and valuations, are best modelled as Bayesian games. While Bayesian modeling is a big step mathematically, in an RL environment it amounts simply to adding a chance node at the root of the game tree. Such a node creates a copy of the game tree for each type combination, which leads to a much larger game tree but, since bidders do not know each others’ types, has a smaller effect on the number of information states. We thus recommend experimenting with bidders that have a small number of types. For example, when we adapted the experiments just described to imperfect-information settings in which each bidder can have one of multiple possible valuations, we have typically observed more interesting, multi-round behavior, including pooling equilibria in which weaker types act like their stronger counterparts to drive better negotiations than they “deserve”.
@@
It is common in auction analysis to model bidders as symmetric. This offers several technical advantages. When all bidders are symmetric, the equilibrium selection problem can mitigated by searching for symmetric equilibria. Furthermore, if all agents behave symmetrically, only a single policy needs to be trained. However, in reality, bidders in spectrum auctions are not usually well-approximated as symmetric, as they often have vastly different resources, business interests, and priorities. Thankfully, such asymmetry is a straightforward addition to a MARL environment, as RL algorithms have little trouble dealing with asymmetric bidders. We advocate for experimenting with varying bidders’ relative strengths.

### v2

@@ Abstract.
Iterative combinatorial auctions are widely used in high stakes settings such as spectrum auctions. Such auctions can be hard to analyze, making it difficult for bidders to determine how to behave and for designers to optimize auction rules to ensure desirable outcomes such as high revenue or welfare. In this paper, we investigate whether multi-agent reinforcement learning (MARL) algorithms can be used to understand iterative combinatorial auctions, given that these algorithms have recently shown empirical success in several other domains. We find that MARL can indeed benefit auction analysis, but that deploying it effectively is nontrivial. We begin by describing modelling decisions that keep the resulting game tractable without sacrificing important features such as imperfect information or asymmetry between bidders. We also discuss how to navigate pitfalls of various MARL algorithms, how to overcome challenges in verifying convergence, and how to generate and interpret multiple equilibria. We illustrate the promise of our resulting approach by using it to evaluate a specific rule change to a clock auction, finding substantially different auction outcomes due to complex changes in bidders’ behavior.
mechanism design, approximate equilibria, multi-agent reinforcement learning
ccs: Applied computing Economicsccs: Theory of computation Exact and approximate computation of equilibriaccs: Theory of computation Multi-agent reinforcement learningjournalyear: 2024copyright: acmlicensedconference: Conference on Economics and Computation; July 8–11, 2024; New Haven, CT, USAbooktitle: Conference on Economics and Computation (EC ’24), July 8–11, 2024, New Haven, CT, USAdoi: 10.1145/3670865.3673644isbn: 979-8-4007-0704-9/24/07
@@ 1. Introduction
Consider the case of spectrum auctions. The iterative combinatorial auctions used to sell radio spectrum can have 50+ page manuals explaining how they evolve over time and which bids are legal to submit (FCC, 2015). They are held too infrequently to learn about from historical data, occurring only once every few years nationally. What’s more, while there are a few popular auction formats, such as Simultaneous Multiple Round Auctions (SMRA) (Milgrom, 2000), Combinatorial Clock Auctions (CCA) (Ausubel et al., 2006), and Clock Auctions [^1] Clock Auctions are combinatorial auctions, but are a distinct auction format from Combinatorial Clock Auctions. (Ausubel, 2004), even within a single format, rules tend to differ from one auction to the next. For example, between the Canadian 700 MHz (ISED, 2012) and 600 MHz (ISED, 2018) auctions, the activity rule was altered; later, between the Canadian 3500 MHz (ISED, 2021) and 3800 MHz (ISED, 2022) auctions, a spectrum cap was added, limiting the maximum a bidder could win in each region. With rules in constant flux, it is important to develop methods for reasoning about the effects of counterfactual policy changes in these auctions. Ideally, we would analyze such problems by finding the Bayes–Nash equilibria of the game induced by each set of auction rules and evaluating how outcomes change in equilibrium. This is the dominant approach taken by theoretical work in economics, and has also been adopted by various computational approaches (e.g. Thompson et al., 2017; Thompson and Leyton-Brown, 2017; Rabinovich et al., 2013). However, there is little hope of identifying Bayes–Nash equilibria of realistic iterative combinatorial auctions, as they are too complex to admit pen-and-paper analysis using known techniques. Thus, the most influential existing work studies highly simplified settings. For instance, for one auction format, Riedel and Wolfstetter (2006) proved that it is an equilibrium for bidders to end the auction in the first round, but their proof applies only to auctions with a single product, infinitesimal price increases, and complete information about bidders’ values. Combinatorial action spaces and multi-round structures imply enormous extensive-form representations, putting them out of reach of traditional equilibrium solvers as well.
@@
An alternative is to simulate the auction under the assumption that bidders play some given, fixed strategy (e.g., Newman et al., 2024; Cary et al., 2007). One can then observe both the effect of a rule change and its sensitivity to changes in bidders’ valuations and to randomness in the auction mechanism. However, combinatorial auctions are rarely strategyproof (see, e.g., Bosshard and Seuken, 2021; Levin and Skrzypacz, 2016). There is thus typically not a singular reasonable strategy to simulate, let alone one with consistent incentive properties across the rules being studied.
@@
Multi-agent reinforcement learning (MARL) offers a middle ground: more computational tractability than classical Bayes–Nash equilibrium computation, but richer strategic reasoning than static simulations. Over the past decade, MARL algorithms have made great strides, achieving superhuman performance in two-player zero-sum games, such as Go (Silver et al., 2016), Shogi (Silver et al., 2017), and Stratego (Pérolat et al., 2022). These games share the property that players do not need to solve the equilibrium selection problem (Harsanyi and Selten, 1988): if a player plays their part of any equilibrium, they will do at least as well against any other opponent equilibrium strategy. Of course, this property does not hold in auctions, nor in zero-sum games having more than two players. It is thus striking that recent work has shown superhuman performance in multiplayer zero-sum games such as poker (Brown and Sandholm, 2019b). Of particular relevance to what follows, Counterfactual Regret Minimization (CFR) (Zinkevich et al., 2007) is a popular, modern MARL algorithm that was recently used to solve Texas Hold ’em poker (Bowling et al., 2015; Brown and Sandholm, 2018). Also impressive was the demonstration of human-level play in the board game Diplomacy (Anthony et al., 2022; Bakhtin et al., 2022), a game sharing other key similarities with our domain: a combinatorial action space and iterative, simultaneous moves.
@@
We foresee no near-term interest from market participants in deploying “superhuman” bots to bid autonomously in complex and high-stakes auctions without human oversight—nobody is going to bet a Fortune 500 company’s future on an RL algorithm running against unknown opponent strategies in a novel economic environment. However, even in such domains, MARL offers promise for evaluating competing mechanisms and gaining economic understanding. Tools from deep learning and reinforcement learning have been used to find revenue-optimal designs for small, single-stage auctions (Dütting et al., 2019, 2023; Nguyen, 2020), optimize reserve prices in repeated auctions (Tang, 2017; Ai et al., 2022), and react to opponent policies in ad auctions (Jin et al., 2018; Banchio and Skrzypacz, 2022). MARL algorithms have also been used to find equilibria in single-round combinatorial auctions (Bosshard et al., 2020; Bichler et al., 2021, 2023) and multi-round dynamic or sequential mechanisms (Thoma et al., 2023a, b; Pieroth et al., 2023). Closest to our precise domain, Pacaud et al. (2023) combine heuristic strategies with Monte-Carlo tree search to find effective bidding strategies in clock auctions; however, they focus on finding strategies that perform well against existing heuristics in auctions with complete information. Beyond auctions, MARL has been used to identify and evaluate complex policy decisions regarding taxation (Zheng et al., 2022), study regulations on platform economies (Wang et al., 2023), and study bargaining dynamics of learning algorithms (Abramowicz, 2020; Li et al., 2023).
@@
This paper aims to use MARL to gain economic insight in (incomplete-information) iterative combinatorial auctions. Such insights could help a human bidder assemble a strategic “playbook” by providing examples of strong bidding both by an agent with their own preferences and by their counterparties; to date, such playbooks are typically derived via pen-and-paper analysis of dramatically simplified subproblems and via observations of expert human play in a handful of “mock auctions”. Insights gleaned from MARL analysis could also help an auction designer to trade off the costs and benefits of candidate rule changes, evaluating economic variables such as revenue, welfare, length of the auction, variance in outcomes, etc. Unfortunately, actually running MARL to understand an iterative combinatorial auction is not nearly as simple as implementing an auction simulator and unleashing multiple copies of some off-the-shelf algorithm. Modeling a spectrum auction verbatim from its rulebook produces games that are vastly too large to solve. One is then immediately faced with important modeling choices, finding ways to reduce the game to a feasible size without abstracting away the most strategically important elements.
@@
Our work makes two main contributions. First, in Section 2, we present a general methodology for studying complex, iterative auctions using MARL. Notably, we recommend simplifying the auction model by abstracting the action space, controlling the number of bidders, and limiting the auction’s length, without giving up on qualitatively important features such as imperfect information or asymmetry between bidders. We describe the landscape of MARL algorithms and important problems that can arise, such as arbitrary behavior when multiple actions lead to the same outcome, and convergence to brittle equilibria that involve unrealistic levels of coordination. We also discuss key challenges in verifying an algorithm’s convergence, and interpreting trained policies in the face of multiple equilibria.
@@
Second, we present a case study showing that, despite the challenges just described, MARL can be used to conduct novel, economically meaningful analysis. We consider a prominent family of iterative combinatorial auctions called clock auctions and, leveraging two modern MARL algorithms, investigate a key design choice: how to process bids when multiple bidders want to reduce demand (submit a drop bid) for the same product to a level below supply. Section 3 describes the problem, how we model it as a game, how we solved it using MARL, and how we assessed performance. Section 4 then goes on to demonstrate that this rule change does indeed give bidders incentive to change their behavior, yielding substantial differences in both revenue and auction length. Furthermore, we show that a designer would incorrectly conclude that the rule change has the opposite effect if they instead modeled bidders as following a straightforward, myopic bidding heuristic.
@@ 2. Methodology: Understanding Iterative Combinatorial Auctions via Multi-Agent Reinforcement Learning
We begin with our methodology for applying MARL to understand iterative combinatorial auctions. First, we discuss how we model an auction, trading off model fidelity with size of the resulting game tree. We then turn to “solving” the game using MARL, discussing the strengths and weaknesses of different MARL algorithm families, the problem of multiple equilibria, and equilibrium refinements (preferring pure equilibria; avoiding brittle equilibria). We conclude by considering how to validate and interpret policies: assessing the extent to which policies have converged to a Bayes–Nash equilibrium and reasoning across multiple equilibria.
@@ 2.1. Modelling an Auction
It is tempting to simply model an auction wholesale, mapping its entire rules manual into a MARL environment, and unleash a MARL algorithm on it. Unfortunately, this approach is unlikely to succeed for iterative combinatorial auctions: modeling many bidders, bundles, and rounds will yield an enormous game tree that is simply infeasible to solve. Much like pen-and-paper analysis, one must abstract away certain elements to buy the ability to accurately model others.
@@
We recommend gaining a solid foundation by starting with simple, easy-to-solve games and scaling up later rather than starting with complex ones that might prove to be intractable. We have been surprised countless times by how even auctions with moderate numbers of bidders, products, and rounds can lead to enormous, unwieldy game trees. However, as we demonstrate in our experiments in Section 3, there is a sweet spot of game sizes out of reach of pen-and-paper analyses, yet tractable for MARL. Here, we describe several high-level decisions that a designer can make to control the size of the game tree. The key metric of importance is usually the number of distinct information states—the number of game states that an agent can distinguish.
@@
Combinatorial auctions typically have enormous action spaces, often corresponding to the set of all legal bundles. The action space determines the branching factor of the game tree, so reining it in has a dramatic effect on game size. One helpful way to limit the size of an action space is to experiment with auctions that have few products or units of each product available, limiting the number of possible bids. However, we caution against reducing the auction down to a single product (e.g., several units of a generic license), as auctions with only one product may not engage with many of the auction’s rules, leaving interesting behavior unstudied. We recommend beginning with auctions offering two or three products.
@@
It is also possible to study auctions with more available bundles by restricting bids, only allowing bidders to choose from a small number of heuristic strategies at each state. For example, these heuristics could include bidding myopically, maintaining the last round’s bid, bidding on an opponent’s holdings, or attempting to drop out of the auction. Care is required to ensure that this restriction does not significantly hamper players’ abilities to best respond to others’ strategies. As a concrete example, it may be tempting to shrink the action space by limiting bidders to only bidding on packages they genuinely value. However, it is well-known that there are auction instances where such simple bidding is never optimal in Bayes–Nash equilibrium (Bosshard and Seuken, 2021). A helpful intermediate option is to allow players to deviate from a given set of heuristics up to $k$ times per auction; this can be useful when it is only strategically important to deviate from known heuristics occasionally, and also can be used to diagnose the effectiveness of the given set of heuristics. This approach has the beneficial side effect of producing strategies that are easier for an analyst to interpret, which can be particularly helpful for developing a playbook. Restricting actions to a smaller set is also likely to make the resulting strategies more useful as meta-strategies in Empirical Game Theory Analysis (Wellman, 2006).
@@
In some auctions, a naive implementation would produce a continuous action space. For instance, both intra-round bidding in clock auctions and the supplementary rounds of CCAs give bidders the ability to report real-valued prices. Continuous action spaces pose a problem for many MARL algorithms, which commonly assume that the action space is discrete. Thus, when possible, we recommend to avoid modelling the part of the auction that admits a continuous action space. When this is impractical, it may be possible to discretize the action space: for instance, allowing bidders to only bid in \$1 increments, or again employing a finite set of heuristics (each of which maps to an unrestricted real value). Alternatively, one can use a MARL algorithm that is designed specifically for continuous action spaces; we caution that these algorithms can make strong assumptions about agents’ policies, requiring bidders to make deterministic bids (Lowe et al., 2017) or to draw their bids from a parametric (e.g., Gaussian) distribution.
@@
A related technical issue is that, in theory, auctions can be arbitrarily long: as long as bidders continue to overdemand products, the auction will continue. These infinitely large game trees do not typically pose a problem to theorists, who can handle them with inductive proofs, nor to reasonable bidders in practice, who would never expose themselves to indefinitely rising prices. However, they do pose a problem to most RL algorithms, which cannot learn to avoid each of these unreasonable bid trajectories without first gaining experience by making them. We recommend constraining agents’ bids to rule out such arbitrary-length bidding sequences. A natural solution is to disallow bidders from making bids that are strictly dominated by dropping out of the auction, ensuring that they are forced to drop out when prices become sufficiently high. Another is modelling bidders as budget-constrained, disallowing bids that would exceed a bidder’s budget if processed. Budgets are a more tenuous modelling choice: while real-world bidders do have finite resources, it may be unrealistic to prohibit a bidder from buying a license that they value more than its cost. We recommend against adding an artificial limit on the length of the auction.222 We experimented with directly limiting the auction’s length, giving all players arbitrary, negative utilities upon hitting this limit. We recommend against this as it can have a dramatic effect on the equilibria of the game. Specifically, we found that weak bidders that would otherwise be allocated nothing could make non-credible threats to bid up to the limit, forcing stronger bidders to concede an item.
@@
In practice, spectrum auctions can involve dozens of bidders. However, many bidders represent small, locally-operating telecoms that have limited ability to make large-scale bids. If smaller bidders must be modeled, consider whether a deterministic heuristic strategy (as opposed to a learned policy) suffices, or whether their action spaces can be restricted to their bundles of interest.
@@
Real auctions, where bidders likely have priors over their opponents’ strategic goals and valuations, are best modeled as Bayesian games. While Bayesian modeling is a big step mathematically, in an RL environment it amounts simply to adding a chance node at the root of the game tree. Such a node creates a copy of the game tree for each type combination, which leads to a much larger game tree but, since bidders do not know each others’ types, has a smaller effect on the number of information states. We thus recommend experimenting with bidders that have a small number of types. For example, when experimented with imperfect-information settings in which each bidder can have one of multiple possible valuations, we have typically observed more interesting, multi-round behavior, than in the perfect-information case, including pooling equilibria in which weaker types act like their stronger counterparts to drive better negotiations than they “deserve”.
@@
It is common in auction analysis to model bidders as symmetric. This offers several technical advantages. When all bidders are symmetric, the equilibrium selection problem can mitigated by searching for symmetric equilibria. Furthermore, if all agents behave symmetrically, only a single policy needs to be trained. However, in reality, bidders in spectrum auctions are not usually well-approximated as symmetric, as they often have vastly different resources, business interests, and priorities. Thankfully, such asymmetry is a straightforward addition to a MARL environment, as RL algorithms have little trouble dealing with asymmetric bidders. We advocate for experimenting with bidders of varying relative strengths.